# A calibrated model for N-body dynamical friction acting on supermassive black holes

### v1

@@ 5.2.2 An eccentric orbit
Figure 12: The median trajectories of $10^{9}\,{\rm M}_{\odot}$ black holes set on an eccentric orbit in a $10^{13}\,{\rm M}_{\odot}$ halo. The median trajectories are computed from 6 random seeds at each resolution level (same as Fig. 11). The arrows indicate the earliest and latest sinking black hole. The black hole dynamics were computed using our correction for the unresolved dynamical friction and the correction for stochasticity. The grey dotted and dashed lines show the sinking time for $M_{\rm BH}/M_{\rm DM}=1000$ predicted by the Chen et al. (2022) and Tremmel et al. (2015) models, respectively.
@@
Figure 13: The trajectory of a $10^{6}\,{\rm M}_{\odot}$ black hole in a $2\times 10^{11}\,{\rm M}_{\odot}$ halo from the HighRes example presented in Tremmel et al. (2015). The red line shows the result of applying our model, while the blue line is the result of the Tremmel et al. (2015) model. The black solid and dashed lines are the analytical and KETJU predictions, respectively.
Finally, we apply our model to the HighRes test case presented in Tremmel et al. (2015). This includes a $2\times 10^{11}\,{\rm M}_{\odot}$ halo with concentration $c=4.5$, in which a black hole with mass $10^{6}\,{\rm M}_{\odot}$ starts off at 2 kpc, with a velocity $v=0.1\,v_{\rm circ}$. The dark matter particle mass is $1.53\times 10^{4}\,{\rm M}_{\odot}$, resulting in $M_{\rm BH}/M_{\rm DM}\approx 60$. The gravitational softening is set to $\epsilon_{g}=77$ pc. In Fig. 13, the red line shows the black hole trajectory using our dynamical friction model, and the blue line shows the trajectory of the black hole using the Tremmel et al. (2015) model. The sinking time prediction for the analytical model and KETJU differs by $\sim$0.05 Gyr. We note that this was the most computationally expensive of our simulations in this work, as the accuracy of the time integration in the analytical model (and consequently the simulations with our model and the model of Tremmel et al. 2015) had to be increased to 0.1 per cent in order to accurately capture the eccentric orbit of the black hole, as confirmed by convergence tests. Our model follows the analytical prediction and that of KETJU almost exactly, while the model of Tremmel et al. (2015) results in the black hole sinking time delayed by almost $0.2$ Gyr.
@@ 6 Conclusions
Of course, our model has only been tested in spherical, equilibrium $\rho\propto r^{-1}$ halos and its validity may change when applied to different density distributions (particularly cored halos, see e.g. Antonini & Merritt 2012), flattened systems and systems undergoing a merger, which would introduce spatial and velocity anisotropy (Peñarrubia et al., 2004). On the other hand, a great advantage of this model and previous models of its kind, is that some dynamical friction is already captured in the simulations, and the correction we apply is only a local one. In other words, the accuracy of the recovered black hole motion should depend only on whether the Chandrasekhar formula represents well the impact of dynamical friction within $\sim\epsilon_{g}$, which, depending on the spatial resolution of the simulation, may be a very small scale. Ultimately, testing the full range of validity of our model requires more sophisticated numerical setups and cosmological zoom simulations, as well as complementary high-resolution simulations with KETJU, all of which we will leave to future work.
@@ Acknowledgements
The authors would like to thank Yueying Ni, Christian Partmann and Simon White for helpful discussions. This research was carried out using the High Performance Computing resources of the FREYA cluster at the Max Planck Computing and Data Facility (MPCDF, https://www.mpcdf.mpg.de) in Garching operated by the Max Planck Society (MPG). This project was developed as part of the Simons Collaboration on “Learning the Universe”.
@@
Appendix A Maximum unresolved impact parameter in the Tremmel model

### v2

@@ 5.2.2 An eccentric orbit
Figure 12: The median trajectories of $10^{9}\,{\rm M}_{\odot}$ black holes set on an eccentric orbit in a $10^{13}\,{\rm M}_{\odot}$ halo. The median trajectories are computed from 6 random seeds at each resolution level (same as Fig. 11). The error bars show the range between the earliest and the latest black hole sinking time. The black hole dynamics were computed using our correction for the unresolved dynamical friction and the correction for stochasticity. The grey dotted and dashed lines show the sinking time for $M_{\rm BH}/M_{\rm DM}=1000$ predicted by the Chen et al. (2022) and Tremmel et al. (2015) models, respectively.
@@
Figure 13: The trajectory of a $10^{6}\,{\rm M}_{\odot}$ black hole in a $2\times 10^{11}\,{\rm M}_{\odot}$ halo from the HighRes example presented in Tremmel et al. (2015). The red line shows the result of applying our model, while the blue line is the result of the Tremmel et al. (2015) model. The black solid and dashed lines are the analytical and KETJU predictions, respectively. All simulations were carried out with $M_{\rm BH}/M_{\rm DM}=376$.
Finally, we apply our model to the initial conditions similar to the HighRes test case presented in Tremmel et al. (2015). This includes a $M_{200}=2\times 10^{11}\,{\rm M}_{\odot}$ halo with concentration $c=4.5$ and $R_{200}=115$ kpc, in which a black hole with mass $10^{6}\,{\rm M}_{\odot}$ starts off at 2 kpc, with a velocity $v=0.1\,v_{\rm circ}$. The dark matter particle mass in our simulations is $2.7\times 10^{3}\,{\rm M}_{\odot}$, resulting in $M_{\rm BH}/M_{\rm DM}\approx 380$. The gravitational softening is set to $\epsilon_{g}=77$ pc, as in Tremmel et al. (2015). In Fig. 13, the red line shows the black hole trajectory using our dynamical friction model, and the blue line shows the trajectory of the black hole using the Tremmel et al. (2015) model. The sinking time prediction for the analytical model and KETJU differs by $\sim$0.05 Gyr. This could be the result of minor stochastic effects we expect for mass ratios $500100<M_{\rm BH}/M_{\rm DM}<$ 500 (see Fig. 4). We also note that this was the most computationally expensive of our simulations in this work, as the accuracy of the time integration in the analytical model (and consequently the simulations with our model and the model of Tremmel et al. 2015) had to be increased to 0.1 per cent in order to accurately capture the eccentric orbit of the black hole, as confirmed by convergence tests. Our model follows the analytical prediction and that of KETJU almost exactly, while the model of Tremmel et al. (2015) results in the black hole sinking time delayed by almost $0.2$ Gyr.
@@ 6 Conclusions
Of course, our model has only been tested in spherical, equilibrium $\rho\propto r^{-1}$ halos and its validity may change when applied to different density distributions (particularly cored halos, see e.g. Antonini & Merritt 2012), flattened systems and systems undergoing a merger, which would introduce spatial and velocity anisotropy (Peñarrubia et al., 2004). On the other hand, a great advantage of this model and previous models of its kind, is that some dynamical friction is already captured in the simulations, and the correction we apply is only a local one. In other words, the accuracy of the recovered black hole motion should depend only on whether the Chandrasekhar formula represents well the impact of dynamical friction within $\sim\epsilon_{g}$, which, depending on the spatial resolution of the simulation, may be a very small scale. Moreover, like models of Tremmel et al. (2015) and Chen et al. (2022), as well as repositioning schemes, our dynamical friction prescription does not conserve momentum. This may lead to unrealistic properties of the particles surrounding the black hole, particularly in regions where the dynamical friction correction is strong. We discuss the implications of this in detail in Appendix C. Ultimately, testing the full range of validity of our model requires more sophisticated numerical setups and cosmological zoom simulations, as well as complementary high-resolution simulations with KETJU, all of which we will leave to future work.
@@ Acknowledgements
The authors would like to thank Yueying Ni, Christian Partmann and Simon White for helpful discussions. This research was carried out using the High Performance Computing resources of the FREYA cluster at the Max Planck Computing and Data Facility (MPCDF, https://www.mpcdf.mpg.de) in Garching operated by the Max Planck Society (MPG). This project was developed as part of the Simons Collaboration on “Learning the Universe”. This research benefited from the use of emcee (Foreman-Mackey et al., 2013), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020) and matplotlib (Hunter, 2007).
@@
Appendix A The diffusion coefficients for a Hernquist halo
In Fig. A1, we demonstrate how the dynamical friction forces compare to the stochastic forces for our fiducial Hernquist halo, with $M_{\rm BH}/M_{\rm DM}=1000$. The diffusion coefficients were computed through explicitly evaluating Eqns. (10), (11) and (12), assuming no softening. It can be seen that the dynamical friction force due to the background particle mass (blue line) is of similar order of magnitude as the second-order diffusion coefficients (red and orange lines), which justifies our approximation of the stochastic force in Eqn. (20). Note that the second-order diffusion coefficient in the perpendicular direction, $D[(\Delta v_{\perp})^{2}]/(2v)$, corresponds to 4 possible directions in which the stochastic force can act, while the parallel second-order diffusion coefficient, $D[(\Delta v_{\parallel})^{2}]/(2v)$, acts only in one direction (the direction of motion of the black hole).
The dashed red lines show how $D[(\Delta v_{\parallel})^{2}]/(2v)$ would change if the mass of the background particles were to increase (reducing $M_{\rm BH}/M_{\rm DM}$), increasing the impact of stochasticity relative to that of dynamical friction.
Appendix B Maximum unresolved impact parameter in the Tremmel model
@@
Appendix C Momentum conservation
Figure A3: Left: fractional change in the magnitude of the total momentum of the simulated system (a $10^{13}\,{\rm M}_{\odot}$ Hernquist halo with a $10^{9}\,{\rm M}_{\odot}$ black hole) when our dynamical friction model is applied. Different colours correspond to different background particle resolution and softening (see Table 1). The black holes sink at $\sim$5 Gyr. Middle: the change in the density profile after 1 Gyr (solid line) and 8 Gyr of evolution (dashed line) for the 4 highest-resolution cases from the left panel. Black lines show the results of the simulation with KETJU (where momentum is essentially conserved, with $\Delta|p|/|p_{0}|\sim 10^{-8}$). The grey dotted line is the analytical Hernquist density profile. The grey shaded band covers the region of the halo containing $10^{9}\,{\rm M}_{\odot}$ of dark matter in the initial conditions, equal to the black hole particle mass. The arrows show the gravitational softening used in each simulation. Right: as the middle panel, but showing the change in 3-dimensional velocity dispersion.
Similarly to the models of Tremmel et al. (2015) and Chen et al. (2022), as well as repositioning schemes, our dynamical friction model does not conserve the momentum of the simulated system. This is demonstrated on the left panel of Fig. A3, where we show the fractional change of the magnitude of momentum compared to the initial conditions. It is evident that in the lower-resolution simulations (where the softening is large), the discrepancy is more significant and the change in momentum can reach $\sim 10$ per cent of the initial value by the time the black hole sinks (at $\sim 5$ Gyr). After the black hole reaches the centre of the halo, the change in the momentum of the system can become extremely high, especially at low resolution. This is because the black hole is not fully stable at the halo centre and, as long as $b_{\rm min}$ is smaller than $b_{\rm max,unres}=\eta\epsilon_{g}$, the deceleration of the black hole in high-density regions computed with our model is extremely large. For the high-resolution case, $M_{\rm BH}/M_{\rm DM}=5000$ and $\epsilon_{g}=0.2$ kpc, since the central regions are well resolved and $b_{\rm min}$ becomes greater than $b_{\rm max,unres}$ (setting the dynamical friction correction to 0) the black hole is able to stay at the centre of the halo and the momentum thereafter is essentially conserved.
As the black hole (or any massive infalling object) is slowed down due to dynamical friction, it loses angular momentum, while the surrounding particles would, in turn, gain momentum. This process has previously been invoked as a mechanism for transforming cold dark matter cusps in the centres of haloes into constant-density cores (El-Zant et al., 2001; El-Zant et al., 2004; Cole et al., 2011). In the middle panel of Fig. A3, we demonstrate that our simulation with KETJU of a $10^{9}\,{\rm M}_{\odot}$ black hole sinking to the centre of a $10^{13}\,{\rm M}_{\odot}$ Hernquist halo (black solid and dashed lines) also predicts a shallower dark matter density profile once the black hole has sunk to the halo centre. An increase in the velocity dispersion is also evident below a radius of 800 pc (where the initial Hernquist profile contains $10^{9}\,{\rm M}_{\odot}$ in dark matter). In coloured lines, we show how these halo properties differ when our dynamical friction model is applied. When the softening is small (< 400 pc), the flattening of the density profile is partially recovered (as momentum exchange still occurs on the scales where dynamical friction is resolved) and the increase in the velocity dispersion of the surrounding particles can also be seen as the surrounding particles are heated by the black hole. When the softening is large (> 400 pc), the slope of the density profile does not substantially change and neither does the velocity dispersion. Although the lower-resolution cases do have a central density core, this is primarily induced by the gravitational softening (Barnes, 2012), while the background particle kinematics remain cold.
In practice, the lack of momentum conservation would lead to underestimated velocity dispersion of background stars and dark matter in the vicinity of the black hole, with the effect being strongest when the gravitational softening is large, for very massive black holes and in dense regions, where the dynamical friction is strong. To circumvent this issue, one could, in principle, artificially inject momentum into the surrounding particles such that total momentum is manifestly conserved. It is, however, unclear how to distribute this momentum, such that the resultant dynamics of the surrounding particles faithfully reflect momentum transfer in dynamical friction. Perhaps, a weighting scheme for particles with distance $d<\eta\epsilon_{g}$ of the black hole that follows Eqn. (9) of Ma et al. (2023) could be used, though this will likely break down at low resolution if the black hole is not stable at the centre of the halo (see the left panel of Fig. A3), leading to an unnecessarily large cumulative momentum injection into the surrounding matter. For now, it is clear that spatial and kinematic properties of particles surrounding a massive black hole in galaxy centres cannot be taken at face value in simulations where momentum exchange due to dynamical friction is not fully accounted for.